# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v2

@@ Appendix C Klich’s determinant formula
$\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)=\text{det}\left(1-\xi e^{A}e^{B}\right)^{-\xi},$ (120)
@@
$\hat{A}=\sum_{i,j=1}^{J}a_{i}^{\dagger}A^{(0)}_{ij}a_{j}+\frac{1}{2}a_{i}^{\dagger}A^{(1)}_{ij}a_{j}^{\dagger}+\frac{1}{2}a_{i}A^{(2)}_{ij}a_{j}=\frac{1}{2}\mathcal{A}-\frac{1}{2}\text{tr}A^{(0)},$ (121)
@@
$\text{Diag}\Lambda=\left(\begin{array}[]{cccccc}\lambda_{1}&\cdots&\lambda_{J}&-\lambda_{J}&\cdots&-\lambda_{1}\end{array}\right),$ (125)
@@
if ($\text{if }\left(ZA-\lambda\right)^{m}\left|\lambda\right>$ $=0$
$\implies\left|\lambda\right>^{T}ZX\left(ZA+\lambda\right)^{m}$ $=0,$ (126)
for $m\geq 1$, the $m\neq 1$ cases correspond to generalized eigenvectors. The specific arrangement of eigenvalues in eq.(125) also imposes a specific structure on the matrix $U$, which is
@@
$\left|\lambda_{i}\right>^{\dagger}\left|\lambda_{j}\right>$ $=\delta_{ij}.$ (129)
@@
$=\left(\begin{array}[]{cccccc}c_{1}^{\bullet}&\cdots&c_{J}^{\bullet}&c_{J}&\cdots&c_{1}\end{array}\right)Z\Lambda\left(\begin{array}[]{cccccc}c_{1}&\cdots&c_{J}&c_{J}^{\bullet}&\cdots&c_{1}^{\bullet}\end{array}\right)^{T},$ (132)
and the commutation relations for the new operators are $[c_{j},c^{\bullet}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $[c_{j},c_{j^{\prime}}]=0$, $[c^{\bullet}_{j},c^{\bullet}_{j^{\prime}}]=0$. Finally,
$\mathcal{A}=\sum_{i=1}^{J}\lambda_{i}\left(c_{i}^{\bullet}c_{i}+c_{i}c_{i}^{\bullet}\right)+\sum_{i,j=1}^{J}N_{ij}c_{i}^{\bullet}c_{j}^{\bullet}+\sum_{i<j,i=1}^{J}R_{ij}c_{i}^{\bullet}c_{j},$ (133)
@@
Define two states $0\right|_{L}$ and $\left|0\right>_{R}$ such that
$0\right|_{L}c^{\bullet}_{i}$ $=0\text{ }\forall i,$ (134)
$c_{i}\left|0\right>_{R}$ $=0\text{ }\forall i.$ (135)
@@
$\left|n_{1},n_{2},\cdots,n_{J}\right>$ $=\frac{{c^{\bullet}_{1}}^{n_{1}}{c^{\bullet}_{2}}^{n_{2}}\cdots{c^{\bullet}_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}\left|0\right>_{R},$ (136)
$n_{1},n_{2},\cdots,n_{J}\right|\right.$ =$0\right|_{L}\frac{{c_{1}}^{n_{1}}{c_{2}}^{n_{2}}\cdots{c_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}.$ (137)
@@
$n_{1},n_{2},\cdots,n_{J}\right|\right.\left|m_{1},m_{2},\cdots,m_{J}\right>=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\cdots\delta_{n_{J},m_{J}},$ (138)
@@
^$\hat{\mathcal{I}}=\sum_{n_{1}=0}^{\infty}\cdots\sum_{n_{J}=0}^{\infty}\left|n_{1},n_{2},\cdots,n_{J}\right>\left<\left<n_{1},n_{2},\cdots,n_{J}\right|\right..$ (139)
@@
$=\prod_{i=1}^{J}\frac{e^{\lambda_{i}/2}}{1-e^{\lambda_{i}}}.$ (140)
The terms $\sum_{i,j=1}^{J}N_{ij}c_{i}^{\bullet}c_{j}^{\bullet}$ and $\sum_{i<j,i=1}^{J}R_{ij}c_{i}^{\bullet}c_{j}$ change the number states in such a way that no combination of them with each other or themselves can contribute to the trace, therefore the only term that contributes is $\sum_{i=1}^{J}\lambda_{i}\left(c_{i}^{\bullet}c_{i}+c_{i}c_{i}^{\bullet}\right)$. This was the reason why we chose a specific arrangement of eigenvalues in eq.(125).
@@
$\text{Tr}e^{\hat{A}}=e^{-\text{tr}A^{(0)}/2}\det(1-UU^{\dagger}Ze^{ZA/2})^{-1},$ (141)
@@
$\left[\text{Tr}e^{\hat{A}}\right]^{2}=\det(Z)e^{-\text{tr}A^{(0)}}\det(1-e^{ZA})^{-1}.$ (142)
We still don’t have an expression which is analogous to the eq.(120), to derive such an expression, consider another operator $\hat{B}$ that can be written in a form similar to eq.(121), then it can be shown that
$\left[\frac{1}{2}\mathcal{A},\frac{1}{2}\mathcal{B}\right]$ $=\frac{1}{2}\mathcal{C},$ (143)
@@
$\left[\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)\right]^{2}=\det(Z)e^{-\text{tr}(A^{(0)}+B^{(0)})}\det(1-e^{ZA}e^{ZB})^{-1}.$ (144)
@@ Appendix D Correlation matrix from Klich’s determinant formula
$=e^{-\text{tr}A^{(0)}}\det\left(\frac{\mathbb{I}_{2J}+Z}{2}+Q^{-1}e^{ZA}Q\frac{\mathbb{I}_{2J}-Z}{2}\right)^{-1},$ (145)
@@
$=$ $\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{\text{Tr}\left(e^{i\mathcal{H}_{>}t}e^{\hat{\chi}}e^{-i\mathcal{H}_{>}t}e^{-\beta\mathcal{H}_{<}}\right)^{2}}{\text{Tr}\left(e^{-\beta\mathcal{H}_{<}}\right)^{2}}}-1}{\delta\chi},$ (146)
@@
$M$ $=\left(\begin{array}[]{cc}M^{(0)}&M^{(1)}\\ M^{(2)}&\left(M^{(0)}\right)^{T}\end{array}\right),$ (149)
@@
$=\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{e^{\pm\delta\chi\text{tr}M^{(0)}}\det\left(1-e^{iZ_{2J}H_{>}t}e^{\delta\chi Z_{2J}M}e^{-iZ_{2J}H_{>}t}e^{-\beta Z_{2J}H_{<}}\right)^{-1}}{\det\left(1-e^{-\beta Z_{2J}H_{<}}\right)^{-1}}}-1}{\delta\chi},$ (150)
where we have used Eq.(144). Expanding the exponentials upto first order in $\delta\chi$, and using the identity $\det\left(\mathbb{I}+\epsilon A\right)\approx 1+\epsilon\text{tr}\left(A\right)+\mathcal{O}\left(\epsilon^{2}\right)$, we get
$F_{>,ij}(t,t)=\lim_{\delta\chi\to 0}\frac{e^{\pm\delta\chi\text{tr}\frac{M^{(0)}}{2}}\left(1+\frac{\delta\chi}{2}\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]\right)^{-1/2}-1}{\delta\chi}.$ (152)
@@
$=-\frac{\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]}{4}.$ (153)
@@
$=-\frac{1}{2}-\frac{\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]}{4}.$ (154)
@@
$=\frac{1}{2}-\frac{\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]}{4}.$ (155)
@@
    $M_{lk}^{(0)}$ =$=\begin{cases}0&l\neq i,k\neq j,\\ 1&l=i,k=j,\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}-\frac{1}{2}\text{tr}M^{(0)}\right),$ (156)
@@
    $M_{lk}^{(1)}$ =$=\begin{cases}1&l=j-J,k=i,\\ 1&l=i,k=j-J,\\ 0&\text{every other element,}\end{cases}$
    $\hat{\chi}$ $=\delta\chi\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}.$ (157)
@@
    $M_{lk}^{(2)}$ =$=\begin{cases}1&l=j,k=i-J,\\ 1&l=i-J,k=j,\\ 0&\text{every other element,}\end{cases}$
    $\hat{\chi}$ $=\delta\chi\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}.$ (158)
@@
    $M_{lk}^{(0)}$ =$=\begin{cases}0&l\neq i-J,k\neq j-J,\\ 1&l=i-J,k=j-J,\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}+\frac{1}{2}\text{tr}M^{(0)}\right).$ (159)

### v4

@@ Appendix C Klich’s determinant formula
$\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)=\text{det}\left(1-\xi e^{A}e^{B}\right)^{-\xi},$ (C.1)
@@
$\hat{A}=\sum_{i,j=1}^{J}a_{i}^{\dagger}A^{(0)}_{ij}a_{j}+\frac{1}{2}a_{i}^{\dagger}A^{(1)}_{ij}a_{j}^{\dagger}+\frac{1}{2}a_{i}A^{(2)}_{ij}a_{j}=\frac{1}{2}\mathcal{A}-\frac{1}{2}\text{tr}A^{(0)},$ (C.2)
@@
$\text{Diag}\Lambda=\left(\begin{array}[]{cccccc}\lambda_{1}&\cdots&\lambda_{J}&-\lambda_{J}&\cdots&-\lambda_{1}\end{array}\right),$ (C.6)
@@
$\text{if }\left(ZA-\lambda\right)^{m}\left|\lambda\right$> $=0$
$\implies\left|\lambda\right>^{T}ZX\left(ZA+\lambda\right)^{m}$ $=0,$ (C.7)
for $m\geq 1$, the $m\neq 1$ cases correspond to generalized eigenvectors. The specific arrangement of eigenvalues in Eq. (C.6) also imposes a specific structure on the matrix $U$, which is
@@
$\left|\lambda_{i}\right>^{\dagger}\left|\lambda_{j}\right$> $=\delta_{ij}.$ (C.10)
@@
$=\left(\begin{array}[]{cccccc}c_{1}^{\bullet}&\cdots&c_{J}^{\bullet}&c_{J}&\cdots&c_{1}\end{array}\right)Z\Lambda\left(\begin{array}[]{cccccc}c_{1}&\cdots&c_{J}&c_{J}^{\bullet}&\cdots&c_{1}^{\bullet}\end{array}\right)^{T},$ (C.13)
and the commutation relations for the new operators are $c_{j},c^{\bullet}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $c_{j},c_{j^{\prime}}]=0$, $c^{\bullet}_{j},c^{\bullet}_{j^{\prime}}]=0$. Finally,
$\mathcal{A}=\sum_{i=1}^{J}\lambda_{i}\left(c_{i}^{\bullet}c_{i}+c_{i}c_{i}^{\bullet}\right)+\sum_{i,j=1}^{J}N_{ij}c_{i}^{\bullet}c_{j}^{\bullet}+\sum_{i<j,i=1}^{J}R_{ij}c_{i}^{\bullet}c_{j},$ (C.14)
@@
Define two states $\left<0\right|_{L}$ and $\left|0\right>_{R}$ such that
$\left<0\right|_{L}c^{\bullet}_{i}$ $=0\text{ }\forall i,$ (C.15)
$c_{i}\left|0\right>_{R}$ $=0\text{ }\forall i.$ (C.16)
@@
$\left|n_{1},n_{2},\cdots,n_{J}\right$> $=\frac{{c^{\bullet}_{1}}^{n_{1}}{c^{\bullet}_{2}}^{n_{2}}\cdots{c^{\bullet}_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}\left|0\right>_{R},$ (C.17)
$\left<\left<n_{1},n_{2},\cdots,n_{J}\right|\right$. $=\left<0\right|_{L}\frac{{c_{1}}^{n_{1}}{c_{2}}^{n_{2}}\cdots{c_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}.$ (C.18)
@@
$\left<\left<n_{1},n_{2},\cdots,n_{J}\right|\right.\left|m_{1},m_{2},\cdots,m_{J}\right>=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\cdots\delta_{n_{J},m_{J}},$ (C.19)
@@
$\hat{\mathcal{I}}=\sum_{n_{1}=0}^{\infty}\cdots\sum_{n_{J}=0}^{\infty}\left|n_{1},n_{2},\cdots,n_{J}\right>\left<\left<n_{1},n_{2},\cdots,n_{J}\right|\right.$. (C.20)
@@
$=\prod_{i=1}^{J}\frac{e^{\lambda_{i}/2}}{1-e^{\lambda_{i}}}.$ (C.21)
The terms $\sum_{i,j=1}^{J}N_{ij}c_{i}^{\bullet}c_{j}^{\bullet}$ and $\sum_{i<j,i=1}^{J}R_{ij}c_{i}^{\bullet}c_{j}$ change the number states in such a way that no combination of them with each other or themselves can contribute to the trace, therefore the only term that contributes is $\sum_{i=1}^{J}\lambda_{i}\left(c_{i}^{\bullet}c_{i}+c_{i}c_{i}^{\bullet}\right)$. This was the reason why we chose a specific arrangement of eigenvalues in Eq. (C.6).
@@
$\text{Tr}e^{\hat{A}}=e^{-\text{tr}A^{(0)}/2}\det(1-UU^{\dagger}Ze^{ZA/2})^{-1},$ (C.22)
@@
$\left[\text{Tr}e^{\hat{A}}\right]^{2}=\det(Z)e^{-\text{tr}A^{(0)}}\det(1-e^{ZA})^{-1}.$ (C.23)
We still don’t have an expression which is analogous to the Eq. (C.1), to derive such an expression, consider another operator $\hat{B}$ that can be written in a form similar to Eq. (C.2), then it can be shown that
$\left[\frac{1}{2}\mathcal{A},\frac{1}{2}\mathcal{B}\right]$ $=\frac{1}{2}\mathcal{C},$ (C.24)
@@
$\left[\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)\right]^{2}=\det(Z)e^{-\text{tr}(A^{(0)}+B^{(0)})}\det(1-e^{ZA}e^{ZB})^{-1}.$ (C.25)
@@ Appendix D Correlation matrix from Klich’s determinant formula
$=e^{-\text{tr}A^{(0)}}\det\left(\frac{\mathbb{I}_{2J}+Z}{2}+Q^{-1}e^{ZA}Q\frac{\mathbb{I}_{2J}-Z}{2}\right)^{-1},$ (D.1)
@@
$=$ $\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{\text{Tr}\left(e^{i\mathcal{H}_{>}t}e^{\hat{\chi}}e^{-i\mathcal{H}_{>}t}e^{-\beta\mathcal{H}_{<}}\right)^{2}}{\text{Tr}\left(e^{-\beta\mathcal{H}_{<}}\right)^{2}}}-1}{\delta\chi},$ (D.2)
@@
$M$ $=\left(\begin{array}[]{cc}M^{(0)}&M^{(1)}\\ M^{(2)}&\left(M^{(0)}\right)^{T}\end{array}\right),$ (D.5)
@@
$=\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{e^{\pm\delta\chi\text{tr}M^{(0)}}\det\left(1-e^{iZ_{2J}H_{>}t}e^{\delta\chi Z_{2J}M}e^{-iZ_{2J}H_{>}t}e^{-\beta Z_{2J}H_{<}}\right)^{-1}}{\det\left(1-e^{-\beta Z_{2J}H_{<}}\right)^{-1}}}-1}{\delta\chi},$ (D.6)
where we have used Eq. (C.25). Expanding the exponentials upto first order in $\delta\chi$, and using the identity $\det\left(\mathbb{I}+\epsilon A\right)\approx 1+\epsilon\text{tr}\left(A\right)+\mathcal{O}\left(\epsilon^{2}\right)$, we get
$F_{>,ij}(t,t)=\lim_{\delta\chi\to 0}\frac{e^{\pm\delta\chi\text{tr}\frac{M^{(0)}}{2}}\left(1+\frac{\delta\chi}{2}\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]\right)^{-1/2}-1}{\delta\chi}.$ (D.8)
@@
$=-\frac{\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]}{4}.$ (D.9)
@@
$=-\frac{1}{2}-\frac{\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]}{4}.$ (D.10)
@@
$=\frac{1}{2}-\frac{\text{tr}\left[Q^{-1}e^{iZH_{>}t}ZMe^{-iZH_{>}t}Q\left(\mathbb{I}_{2J}-Z\right)\right]}{4}.$ (D.11)
@@
    $M_{lk}^{(0)}$ $=\begin{cases}0&l\neq i,k\neq j,\\ 1&l=i,k=j,\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}-\frac{1}{2}\text{tr}M^{(0)}\right),$ (D.12)
@@
    $M_{lk}^{(1)}$ $=\begin{cases}1&l=j-J,k=i,\\ 1&l=i,k=j-J,\\ 0&\text{every other element,}\end{cases}$
    $\hat{\chi}$ $=\delta\chi\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}.$ (D.13)
@@
    $M_{lk}^{(2)}$ $=\begin{cases}1&l=j,k=i-J,\\ 1&l=i-J,k=j,\\ 0&\text{every other element,}\end{cases}$
    $\hat{\chi}$ $=\delta\chi\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}.$ (D.14)
@@
    $M_{lk}^{(0)}$ $=\begin{cases}0&l\neq i-J,k\neq j-J,\\ 1&l=i-J,k=j-J,\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}+\frac{1}{2}\text{tr}M^{(0)}\right).$ (D.15)
@@
Appendix E Hawking-Unruh radiation on a lattice
In this appendix we argue that the Hawking radiation can indeed be obtained on a lattice if the change in the parameters of the system, equivalently the change in metric, is smooth enough (with some restrictions). However as also pointed out in the main text, there will be essential differences between continuum and lattice models. To begin let us consider the following Lagrangian
$\mathcal{L}=\int dx\left(\frac{1}{2u}\dot{\phi}^{2}-\frac{v\left(x\right)}{u}\dot{\phi}\partial_{x}\phi+\frac{v^{2}\left(x\right)-u^{2}}{2u}\left(\partial_{x}\phi\right)^{2}\right)\ .$ (E.1)
This form of Lagrangian is what most works on analog horizons try to emulate [43]. An important aspect of this field theory is that the determinant of the metric remains constant. This is in contrast to our proposal, where we only modulate the inductance as a function of position, which does not conserve the determinant of the metric in the continuum limit. But our goal here is to compare the result of lattice realisations and continuum realisations, where we choose to consider the simplest possible case with constant determinant.
Now we set this field theory on a one dimensional infinite lattice. The discretised version of the Lagrangian is
$\mathcal{L}=\frac{\Delta x}{u}\sum_{j}\dot{\phi}_{j}^{2}-\sum_{j}\frac{v_{j}}{u}\dot{\phi}_{j}\left(\phi_{j+1}-\phi_{j-1}\right)-\sum_{j}\frac{u^{2}-v_{j}^{2}}{4u\Delta x}\left(\phi_{j+1}-\phi_{j-1}\right)^{2}\ .$ (E.2)
Note that we consider next-to-nearest neighbour coupling in order to avoid any issues due to instabilities unrelated to the horizon (see also main text). Therefore, in this model the lattice constant is best defined as $2\Delta x$. The conjugate field is
$\pi_{j}=\frac{1}{2\Delta x}\frac{\partial\mathcal{L}}{\partial\dot{\phi}_{j}}=\frac{1}{u}\dot{\phi}_{j}-\frac{v_{j}}{2u\Delta x}\left(\phi_{j+1}-\phi_{j-1}\right)\ .$ (E.3)
Using the Legendre transform we can write the Hamiltonian
$\mathcal{H}=2\Delta x\frac{u}{2}\sum_{j}\pi_{j}^{2}+\frac{v_{j}}{2}\sum_{j}\left\{\pi_{j},\left(\phi_{j+1}-\phi_{j-1}\right)\right\}+\sum_{j}\frac{u}{4\Delta x}\left(\phi_{j+1}-\phi_{j-1}\right)^{2}\ .$ (E.4)
Now the equations of motion will be
$2\Delta x\ddot{\phi}_{j}=$ $v_{j+1}\dot{\phi}_{j+1}-v_{j-1}\dot{\phi}_{j-1}+v_{j}\left(\dot{\phi}_{j+1}-\dot{\phi}_{j-1}\right)$
$-\frac{u^{2}-v_{j-1}^{2}}{2\Delta x}\left(\phi_{j}-\phi_{j-2}\right)-\frac{u^{2}-v_{j+1}^{2}}{2\Delta x}\left(\phi_{j}-\phi_{j+2}\right)\ .$ (E.5)
For the translationally invariant case ($v_{j}=v$) this reduces to
$2\Delta x\ddot{\phi}_{j}=2v\left(\dot{\phi}_{j+1}-\dot{\phi}_{j-1}\right)-\frac{u^{2}-v^{2}}{2\Delta x}\left(2\phi_{j}-\phi_{j-2}-\phi_{j+2}\right)\ .$ (E.6)
Using the following ansatz $\phi_{j}(t)=e^{i\omega t}e^{ik\Delta xj}$, we get the dispersion relation
$\omega_{\pm}(k)$ $=v\frac{\sin\left(k\Delta x\right)}{\Delta x}\pm u\left|\frac{\sin\left(k\Delta x\right)}{\Delta x}\right|\ .$ (E.7)
In the limit $\Delta x\rightarrow 0$, this yields the same dispersion relation as in the continuum case. Additionally, in this dispersion relation the sector near $k=0$ and $k=\pi$ mirror each other which is reminiscent of the dispersion relation in Eq. (3).
We now come to the inhomogeneous system. Since the Lagrangian is time-independent we can still write the ansatz $\phi_{j}\left(t\right)=e^{i\omega t}\phi_{j}$, which gives us the following recurrence relation
$-\omega^{2}\frac{2\Delta x}{u}\phi_{j}=$ $i\omega\frac{v_{j+1}}{u}\phi_{j+1}-i\omega\frac{v_{j-1}}{u}\phi_{j-1}+i\omega\frac{v_{j}}{u}\left(\phi_{j+1}-\phi_{j-1}\right)$
$-\frac{u^{2}-v_{j-1}^{2}}{2u\Delta x}\left(\phi_{j}-\phi_{j-2}\right)-\frac{u^{2}-v_{j+1}^{2}}{2u\Delta x}\left(\phi_{j}-\phi_{j+2}\right)\ ,$ (E.8)
which we can re-write as
$\phi_{j+2}=$ $\phi_{j}-\frac{4\omega^{2}\Delta x^{2}}{u^{2}-v_{j+1}^{2}}\phi_{j}+\frac{2i\omega\Delta x}{u^{2}-v_{j+1}^{2}}\left(v_{j-1}\phi_{j-1}-v_{j+1}\phi_{j+1}\right)$
$-\frac{2i\omega\Delta x}{u^{2}-v_{j+1}^{2}}v_{j}\left(\phi_{j+1}-\phi_{j-1}\right)+\frac{u^{2}-v_{j-1}^{2}}{u^{2}-v_{j+1}^{2}}\left(\phi_{j}-\phi_{j-2}\right)\ .$ (E.9)
We are interested in the solution to this recurrence relation for the following profile
$v_{j}=\begin{cases}v_{0}&j\to-\infty,\\ v_{1}&j\to\infty,\end{cases}$ (E.10)
such that $v_{0}<u<v_{1}$. With this profile of the parameter $v_{j}$, in the asymptotically left region the signals will move in both left and right directions while in asymptotically right region the signals will only move towards right, and somewhere in the boundary between these two region for some $j^{\prime}$ we will have $v_{j^{\prime}}\leq u<v_{j^{\prime}+1}$, this is the event horizon.
In general, solving the recurrence relation in Eq. (E) for an arbitrary profile of $v_{j}$ is not analytically possible, therefore we will tackle this problem numerically. We use a linear profile for the parameter $v_{j}$ (for a smooth horizon, we can always approximate the change in its neighbourhood as linear) such that
$v_{j}=\begin{cases}0&j<-J,\\ v_{\text{max}}\frac{j+J}{2J}&-J\leq j\leq J,\\ v_{\text{max}}&j>J,\end{cases}$ (E.11)
where $v_{\text{max}}=\frac{2u}{1-\frac{\alpha}{J}}$ and $\alpha$ (with $0<\alpha<1$, hence $\alpha\ll J$) is a real number that decides a particular detail of the discretisation of $v_{j}$ with respect to $u$. For example, if $\alpha=0$ then $v_{j=0}=u$, or if $\alpha=0.5$ then $u$ will lie exactly in middle of $v_{j^{\prime}}$ and $v_{j^{\prime}+1}$. That is, $\alpha$ controls whether the profile $v_{j}$ actually “hits” the velocity $u$ or not. In fact, for $\alpha=0,1$, the problem is ill-defined as the above recurrence relation contains singularities. For a well-defined answer, the parameter $\alpha$ should lie between 0 and 1 (as specified above).
To construct scattering modes for the Hamiltonian we start with the following incoming plane wave
$\phi_{\text{in}}=\sum_{j<-J}e^{ik\Delta xj}\quad 0<k\Delta x\ll 1,$ (E.12)
in the region with $v_{j}=0$ and energy $\omega_{+}(k)=u\sin\left(k\Delta x\right)/\Delta x$. Using the recurrence relation we obtain the transmitted mode as shown in Fig. 8, from this we conclude that the scattering modes with positive energy are of the form
$\phi_{k}=\sum_{j<-J}e^{ik\Delta xj}+C(\omega_{k,+})\sum_{j\geq-J}\left(e^{i\widetilde{k}\Delta xj}+e^{i\left(\pi-\widetilde{k}\Delta x\right)j}\right),$ (E.13)
where
$\widetilde{k}$ $=\frac{1}{\Delta x}\arcsin{\frac{\omega_{k,+}\Delta x}{v_{\text{max}}-u}},$ (E.14)
is obtained by energy matching and $C(\omega_{k,+})$ is a constant depending on the energy of the incoming wave. To extract the form of the constant $C(\omega_{k,+})$, we use the plot in Fig. 9 (a) and find that
$C(\omega_{k,+})=\gamma e^{\frac{\pi}{a}\omega_{k,+}}\ ,$ (E.15)
with fitting parameters $\gamma$ and $a$.
Figure 8: In this figure we have plotted the absolute value of $\phi_{j}$ as a function of $j$. Since the incoming mode was chosen to be a plane wave therefore we see $\left|\phi_{j}\right|=1$ for j<Jj<-J. The transmitted mode (orange dotted lines), obtained numerically, shows a beating pattern. To confirm that the beating pattern is produced by a wave vector near $\widetilde{k}=0$ and the other one near $\widetilde{k}=\pi$, we have also plotted a function $A\left(e^{i\widetilde{k}\Delta xj}+e^{i\left(\pi-\widetilde{k}\Delta x\right)j)}\right)$ on the same figure. We see that the theoretical and numerical transmitted modes agree with each other if we scale the theoretical one appropriately using the constant AA.
The factor $e^{\pi\omega_{k,+}/a}$ is the expected Hawking factor where $a$ is the analog surface gravity. Indeed, from Fig. 9 (b) we can see that it is given by the derivative of the group velocity across the horizon,
$a=\frac{v_{\text{max}}}{2J\Delta x},$ (E.16)
i.e., it is the discrete derivative of the group velocity. In the continuum version of this problem, $\gamma$ should be $1$. In the here studied lattice model however, the $\gamma$ term is nontrivial. As alluded to in the main text, $\gamma$ depends on the exact details of the discretisation of $v_{j}$, which we discussed above. If $\alpha=0.5$ and therefore $u$ lies exactly in the middle of some $v_{j^{\prime}}$ and $v_{j^{\prime}+1}$ then $\gamma=1$. As we change $\alpha$ and therefore the asymmetry of $u$ with respect to $v_{j^{\prime}}$ and $v_{j^{\prime}+1}$, the value of $\gamma$ diverges as we approach the singular case $\alpha=0$ (see Fig. 9 (b)). This can also be seen from the recurrence relation in Eq. (E), where if some $v_{j}=u$, the equation becomes non-analytic. Another important feature we can gleam from Eq. (E.13) is that the incoming mode scatters into two different modes with same energy, but with different wave vector: one near $\widetilde{k}=0$ and another near $\widetilde{k}=\pi$. This in turn implies that even for the case of smooth variation of parameters, in a lattice model, the ambiguity between black/white hole nature of the horizon still remains.
[…]
Figure 9: a) In this figure we have plotted the logarithm of maximum of $\left|\phi_{j}(k)\right|$, where $k$ is the wave vector of the incoming mode, against $k$ to extract the constant $C(\omega_{k,+})$. Note that only the transmitted mode has been plotted. We see that for different values of $\alpha$ the log plots differ by their intercepts, this is the factor $\gamma$ that depends on the details of discretisation of $v_{j}$. b) Here we have isolated the plot of $\alpha=0.5$, such that $\gamma=1.0$, to confirm that our theoretical prediction of the surface gravity $a=v_{\text{max}}/2J\Delta x$ matches the numerical calculation, which we can see it indeed does.